OpenStudy (anonymous):

Use synthetic division to factorize this x^3 + 3x^2 -4. Show working out thanks.

Directrix (directrix):

Do you know the Rational Root Theorem? I think we will need it.

OpenStudy (anonymous):

not quite, but can u just me the working out, as long as u show me the working out I should understand.

Directrix (directrix):

The rational root theorem generates a list of possible rational roots. From that list, we choose one root to test using synthetic division. No list means there are no divisiors to test. If you are given a divisor, then synthetic division can be used to work out a quotient. But, here we have no way to get a divisor based on your current knowledge. So, I do not know how to approach the problem.

OpenStudy (anonymous):

well I did figure out that one root is (x-1) and I did do the synthetic division, but I just got the wrong answer for some reason.

Directrix (directrix):

If (x-1) is a factor, then x = 1 is a root. I'll run x = 1 through. First, I have to get the coefficents. x^3 + 3x^2 -4 = 1* x^2 + 3*x^2 + 0*x - 4

OpenStudy (callisto):

By factor theorem, f(1) = 1^3 + 3(1)^2 -4 =0  x^3 + 3x^2 -4 = (x-1)(x^2+4x+4) = (x-1)(x+2)^2